### v1

# Accelerating Molecular Dynamics Simulations with Foundation Neural Network Models using Multiple Time-Step and Distillation

Abstract
We present a strategy to accelerate molecular dynamics simulations using foundation neural network models. To do so, we apply a dual-level neural network multi-time-step (MTS) strategy where the target accurate potential is coupled to a simpler but faster model obtained via a distillation process. Thus, the 3.5 Å-cutoff distilled model is sufficient to capture the fast varying forces, i.e. mainly bonded interactions, from the accurate potential allowing its use in a reversible reference system propagator algorithms (RESPA)-like formalism. The approach conserves accuracy, preserving both static and dynamical properties, while enabling to evaluate the costly model only every 3 to 6 fs depending on the system. Consequently, large simulation speedups over standard 1 fs integration are observed: 4-fold in homogeneous systems and 2.3-fold in large solvated proteins. Such a strategy is applicable to any neural network potential and reduces their performance gap with classical force fields.
1 Introduction
Neural Network potentials (NNPs) 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15 have emerged as powerful tools to perform molecular dynamics (MD) simulations, offering near-quantum mechanical accuracy at a fraction of the computational cost of ab initio methods. Presently, such models can be generalized up to Foundation Models covering complete area of applications of molecular dynamics ranging from Material Science to Chemistry, Biology and Drug Design 16, 17, 18, 19, 20. By learning the potential energy surface from reference quantum chemistry data, these models enable simulations of large and complex molecular systems with significantly improved fidelity compared to classical force fields21. However, this increase in accuracy comes at a cost: NNPs are substantially more expensive to evaluate than traditional empirical potentials. This limits their applicability in term of system sizes and long-timescale simulations.
One of the primary bottlenecks of performing MD with ML potentials lies in the time integration itself. The equations of motion must be solved with a small time step - typically in the order of o emtosecond - to resolve high-frequency motions such as bond vibrations. With computationally intensive ML models, this results in a high number of expensive force evaluations, further amplifying the overall simulation cost.
Multi-Time-Step (MTS) integrators 22, 23, 24 offer a well-established strategy to address this challenge. Originally introduced in the context of classical molecular simulations, MTS methods such as RESPA (Reference System Propagator Algorithm) 22 exploit the separation of timescales between different components of the forces to reduce the number of expensive evaluations. By integrating fast-changing forces with a small time step and slow-changing ones with a larger time step, significant performance gains can be achieved without compromising accuracy.
Despite their proven utility in classical force fields, MTS schemes have not yet been adapted to the context of ML-based force field to the best of our knowledge. Indeed, in this context force decomposition is not naturally given by physical interactions. This represents an important opportunity: the use of multiple neural networks of differing complexity and inference cost could enable efficient MTS schemes specifically tailored to ML potentials.
In this work, we propose the first implementation of a RESPA-based MTS scheme for ML potentials. We integrate this approach into the FeNNol library 25, a neural network framework for molecular simulations that can be coupled to the scalable Deep-HP machine learning interface 26 present in the GPU-accelerated Tinker-HP molecular dynamics package 27, 28. Our method associates a small, fast NNP to the FeNNix-Bio1(M) foundation model 19, leveraging the speed of the former to reduce the frequency of the expensive evaluations of the latter. We systematically evaluate the method using various systems including water solvent, solvated small molecules and proteins in the condensed phase, exploring different hyperparameter settings and model sizes in order to design the fastest setup while maintaining accuracy. We also assess the use of a foundation model as a general-purpose fast component to improve transferability across system.
Our results show that this hybrid RESPA-like scheme enables substantial computational speed-ups with limited loss of accuracy, offering a practical route to scalable and efficient Molecular Dynamics with foundation machine learning models and NNPs in general.
2 Methods
2.1 Integration scheme
In this work, we use two neural network potentials of differing complexity to implement a MTS integration scheme. In this scheme, the dynamics of a cheaper NNP is integrated with a small time step (1 fs) and is periodically corrected (every 2 to 6 steps depending on the system) by the force difference between the small NNP and a larger reference model, here the FeNNix-Bio1(M) model 19. This enables to recover the dynamics of the larger NNP without evaluating its forces at every time step, thus improving the computational efficiency. In particular, we employ the BAOAB-RESPA integration scheme 24 outlined in Algorithm 1, that we implemented in the FeNNol library. For simulations in Tinker-HP, we use the implementation described in 24 and the Deep-HP interface 26 for calling models.
In Algorithm 1, denotes the system’s coordinates, its velocities, the outer time step, the mass and the number of inner steps with time step . denotes the reference FeNNix-Bio1(M) machine-learned force field evaluated at configuration and denotes the cheaper model. We describe the neural networks architectures and the distillation strategies employed to train the model in the next sections.
2.2 Neural Network Architectures
The reference model that we intend to accelerate corresponds to the FeNNix-Bio1(M) model, trained on a broad and diverse dataset. This model is based on a range-separated equivariant transformer architecture where close-range and long-range interactions are described with different spatial resolutions and dedicated attention heads. Its receptive field is in total, via two message-passing interactions. Details about the full architecture are provided in ref. 19. The second, lighter-weight, model that is used in inner steps of the RESPA scheme uses the same base architecture but with reduced capacity and removes the long-range attention heads in order to enable faster inference and lower computational cost. Its receptive field is only (one message-passing interaction), making it much more short-sighted and allowing it to focus mostly on fast-varying ”bonded” forces. The model hyperparameters that we used in the following numerical experiments are provided in Supporting Information.
2.3 Training procedures
We derive the model from the reference one via a knowledge distillation procedure 29, 30 where the distilled model is trained on data labeled with the FeNNix-Bio1(M) model instead of DFT. This ensures that the forces used in the inner loop of the RESPA scheme are as close to the reference ones as possible, minimizing the correction necessary to recover the correct dynamics and thus the frequency of its application. In practice, the small model comes in two flavors, either i) as an on-the-fly system-specific model or ii) as a generic model. We briefly describe here these two distillation strategies. Specific details about the training parameters can be found in Supporting Information.
System-specific model
For each system, a reference dataset is generated by running a short MD simulation (less than a nanosecond) using the reference model. For proteins and, more generally, for large systems, a fragmentation strategy similar to the one proposed in ref. 31 was employed to reduce the computational burden while retaining local structural information. We then evaluate the energies and forces of the collected frames with the reference model. The on-the-fly system-specific model is then trained on this dataset, resulting in a model about 10 times faster than FeNNix-Bio1(M) with our current setup.
Generic model
In addition to system-specific distilled models, we also propose a generic fast model trained on a chemically diverse dataset, enabling broader applicability and faster deployment in new systems. To construct this transferable potential, we generate a training set by evaluating FeNNix-Bio1(M) on a subset of conformations from the SPICE2 dataset 32, 33, which contains a wide variety of small organic molecules and biologically-relevant complexes. The generic model is trained on the obtained dataset, resulting in a transferable model that captures general chemical patterns and can be reused across systems. This generic model can be used directly in the inner loop of our MTS scheme, or serve as an initialization point for further fine-tuning on a target system. This offers a compromise between generality and accuracy, especially when system-specific data is limited or a rapid deployment is desired.
3 Numerical experiments
The results obtained with the MTS scheme for each of the following experiments are compared to a single time step (STS) integrator using the FeNNix-Bio1(M) force field with a time step of . The STS scheme employs a BAOAB Langevin integrator 34, 35.
3.1 Bulk Water
First, to investigate the robustness of our models and integrator, we performed a series of stability tests on a small water box containing 648 atoms. Simulations were carried out using the system-specific model and the generic model, with external time steps ranging from to . For each configuration, we monitored dynamical and thermodynamical observables, including diffusion coefficients, kinetic and potential energies, and temperature. Diffusion coefficients were computed using Tinker36 with the Einstein formula. The stability of the trajectory was evaluated over the full course of the simulation, with qualitative indicators reported in table 1. Additional tests were performed with hydrogen mass repartitioning (HMR), enabling larger time steps.
| Diffusion | Temperature | |||
| System | Generic | System | Generic | |
| STS | 2.2 | 300.19.6 | ||
| 2.2 | 2.1 | 300.59.6 | 300.29.6 | |
| 2.3 | 2.6 | 315.210.6 | 333.312.2 | |
| HMR | 2.2 | 2.4 | 300.89.5 | 303.49.9 |
| HMR | 2.3 | 2.5 | 302.79.8 | 304.110.2 |
| HMR | 2.5 | 2.0 | 304.59.7 | 305.99.9 |
| HMR | NaN | 3.3 | NaN | 327.511.1 |
At short external time step (), both specific and generic model produced stable trajectories with diffusion coefficients and temperatures close to the reference STS values. At , simulations remained stable, though the generic and system-specific model displayed a marked increase in both diffusion coefficient and temperature, suggesting the onset of integration artifacts.
The introduction of HMR enables stable simulations up to , with dynamical and thermodynamical properties remaining within reasonable agreement between STS and MTS integrator for both models. However, at and beyond, instabilities systematically occurred, reflected by divergent values in Table 1 or nonphysical increase in diffusion and temperature. These results highlight that HMR provides a substantial extension of the accessible time step, effectively doubling the stability range compared to standard masses. Nevertheless, care must be taken to avoid excessive increases as instabilities rapidly emerge beyond . In practice, time steps of to without HMR, and up to 5- with HMR, appear to offer the most reliable compromise between computational efficiency, stability, and physical accuracy.
It is well documented that limits in the larger time step usable in the context of MTS schemes in molecular dynamics are related to the coupling of this largest time step with the highest-frequency vibrational modes involved in the system: the so-called resonance effects 37, 38, 39, 40, 41, 42, 43, 44. A simple diagnostic can be made to assess these aspects by considering the velocity autocorrelation spectra in this setup. Figure 1 illustrates this in the context of bulk water where we see the artifacts produced by the periodic MTS force correction (annotated in the inset of Figure 1) progressively coupling with the overtone of the O-H stretching mode (around or with HMR) as the outer time step increases. This diagnostic confirms the limits of the outer time step to with hydrogen mass repartitioning and - without.
We further assess the robustness of our approach by computing the radial distribution function of bulk water. As shown in Figure 2, the MTS simulations (with both generic and system-specific models) with an outer time step of and HMR correctly reproduce the STS results within statistical uncertainties. Overall, with this setup, we obtain a speedup of around 4 compared to STS with time step for bulk water simulations.
3.2 Solvated Molecules
To further assess the accuracy of the proposed MTS scheme, we turn to small solvated molecules. We observe stability limits similar to those for bulk water for the five molecules that we tested: ethanol, benzene, trimethylamine, diethylsulfide, and acetic acid. We then evaluated the hydration free energies (HFEs) of these (as well as water). Calculations were performed leveraging the alchemical Lambda-ABF 45 method, employing both system-specific models trained on-the-fly and a generic small foundation model. All simulations were carried out using the same MTS integration scheme described above, with an inner time step of and an outer time step of .
The predicted HFEs were compared to the STS reference values, as shown in Figure 3. The system-specific model achieved a mean absolute error (MAE) of , a root-mean-square error (RMSE) of , and a coefficient of determination () of . The generic foundation model also showed good performance, with an MAE of , an RMSE of , and of .
These results demonstrate that the MTS scheme preserves high accuracy in free energy calculations while benefiting from a significant reduction in computational cost. They also show that both system-specific and generic models are capable of closely reproducing STS-level results.
3.3 Protein-ligand complexes
To validate our scheme on a biologically relevant system, we performed a NVT molecular dynamics simulation of the lysozyme–phenol complex (PDB ID: 4I7L) in explicit water. We first resorted to a setup of outer time step and HMR which is stable for bulk water, but this led to instabilities after of simulation. Figure 4(a) shows the velocity autocorrelation spectra computed over a short simulation with both STS and MTS with HMR enabled. The close similarity of the obtained spectra indicates that the instabilities are not due to the resonance phenomena described above.
In practice, this source of instability lies in the presence of ”holes” in the potential energy surface of the small model, yielding large differences between its forces and the target ones. This is illustrated in Figure 4(b) showing the distribution of the norm of the difference of forces between FeNNix-Bio1(M) and the distilled model (either finetuned on system-specific fragments or the generic one) along an STS trajectory of the protein-ligand system. We observe a dense distribution of values close to and a long tailed distribution starting from . The latter is associated to large nonphysical force discrepancies yielding the infrequent instabilities that we observe in the MTS simulation.
We attempted to mitigate this issue by clipping the force difference to exclude unrealistic values while preserving the meaningful ones (in practice, we used a threshold of on force components). While improving the stability, this setup still leads to the rare occurrence of crashes due to large differences when the time step is equal to or larger than . We expect that an active learning loop where the ”holes” are first identified and then included in a fine-tuning stage will systematically improve the stability of our MTS scheme with larger time steps; this will be the focus of a future work.
b) Distribution (in log-scale) of the norm of the force differences for the generic model, the fine-tuned model trained on forces, energies, and atomic energies, and the fine-tuned model trained on forces and energy, with respect to the FENNIX-Bio1(M) reference model. The distributions are estimated over the first 500 frames of a phenol–lysozyme in water simulation.
Using an inner time step of and an outer time step of with HMR, stable simulations over were obtained while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5. This setup provides a speedup of 2.3 with respect to the STS simulation with time step, strongly accelerating production simulations with biological systems.
4 Conclusion
We have introduced a practical and efficient Multiple Time Step (MTS) scheme tailored to machine-learned force fields, enabling significant acceleration of molecular dynamics simulations without compromising accuracy for both static and dynamical properties. Our approach, based on a RESPA-like scheme, couples a fast, distilled neural network model with the accurate FeNNix-Bio1(M) reference potential. We demonstrated two complementary distillation strategies: (i) on-the-fly system-specific model distillation, and (ii) the use of a lightweight transferable model for broader applicability.
We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.3 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties. Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently. All in all, our approach allows to reach nearly of simulation per day on a single A100 GPU for a realistic protein-ligand complex while preserving the ab initio-like accuracy of the FeNNix-Bio1(M) model. In combination with well-established accelerated sampling schemes 47, 48, 49, this enables truly large-scale simulations with foundation neural neural network potentials.
Future work will focus on expanding the applicability of this approach through two main avenues. First, randomized time stepping could be explored to further increase the effective time step, while allowing for controlled velocity rescaling as proposed in stochastic RESPA variants, i.e. JUMP integrators 44. Second, system-specific models can be fine-tuned more efficiently using fragment-based active-learning strategies to progressively fill-out ”holes” in the small model’s potential energy surface, potentially improving stability with even larger time steps for more complex chemical systems.
Acknowledgments
This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). Computations have been performed at IDRIS (Jean Zay) on GENCI Grants: no A0150712052 (J.-P. P.).

### v2

@@
 The approach conserves accuracy, preserving both static and dynamical properties, while enabling to evaluate the costly model only every 3 to 6 fs depending on the system.
-Consequently, large simulation speedups over standard 1 fs integration are observed: 4-fold in homogeneous systems and 2.3-fold in large solvated proteins.
+Consequently, large simulation speedups over standard 1 fs integration are observed: 4-fold in homogeneous systems and 2.7-fold in large solvated proteins.
 Such a strategy is applicable to any neural network potential and reduces their performance gap with classical force fields.
@@
 Calculations were performed leveraging the alchemical Lambda-ABF 45 method, employing both system-specific models trained on-the-fly and a generic small foundation model.
-All simulations were carried out using the same MTS integration scheme described above, with an inner time step of and an outer time step of .
-The predicted HFEs were compared to the STS reference values, as shown in Figure 3.
+All simulations were carried out using the same MTS integration scheme described above, with an inner time step of and an outer time step of and HMR.
+For benzene MTS simulations using the generic model, we needed to reduce the outer time step to to avoid instabilities.
+This confirms that the system specific model is more robust than the generic one.
+Figure 3 shows the predicted HFEs compared to the STS reference values.
 The system-specific model achieved a mean absolute error (MAE) of , a root-mean-square error (RMSE) of , and a coefficient of determination () of .
@@
 The distributions are estimated over the first 500 frames of a phenol–lysozyme in water simulation.
-Using an inner time step of and an outer time step of with HMR, stable simulations over were obtained while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5.
-This setup provides a speedup of 2.3 with respect to the STS simulation with time step, strongly accelerating production simulations with biological systems.
+Using an inner time step of and an outer time step of with HMR, we obtain stable simulations over while preserving the protein’s structure and the ligand’s binding mode as shown in Figure 5.
+This setup provides a speedup of 2.7 with respect to the STS simulation with time step, strongly accelerating production simulations with biological systems.
 4 Conclusion
@@
 We demonstrated two complementary distillation strategies: (i) on-the-fly system-specific model distillation, and (ii) the use of a lightweight transferable model for broader applicability.
-We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.3 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties.
+We achieve substantial speedups – up to 4 in mostly homogeneous systems (bulk water and solvated small molecules) and 2.7 in large protein-ligand complexes – while preserving key physical observables such as radial distribution functions, hydration free energies, diffusion coefficients, and protein-ligand structural properties.
 Note that this is a first estimation of the computational gains as the code has yet to be optimized further to handle the dual-level approach more efficiently.